# MusicHiFi: Fast High-Fidelity Stereo Vocoding

### v2

@@ Abstract
Diffusion-based audio and music generation models commonly generate music by constructing an image representation of audio (e.g., a mel-spectrogram) and then converting it to audio using a phase reconstruction model or vocoder. Typical vocoders, however, produce monophonic audio at lower resolutions (e.g., 16-24 kHz), which limits their effectiveness. We propose MusicHiFi — an efficient high-fidelity stereophonic vocoder. Our method employs a cascade of three generative adversarial networks (GANs) that convert low-resolution mel-spectrograms to audio, upsamples to high-resolution audio via bandwidth expansion, and upmixes to stereophonic audio. Compared to previous work, we propose 1) a unified GAN-based generator and discriminator architecture and training procedure for each stage of our cascade, 2) a new fast, near downsampling-compatible bandwidth extension module, and 3) a new fast downmix-compatible mono-to-stereo upmixer that ensures the preservation of monophonic content in the output. We evaluate our approach using both objective and subjective listening tests and find our approach yields comparable or better audio quality, better spatialization control, and significantly faster inference speed compared to past work. Sound examples are at https://MusicHiFi.github.io/web/.
@@ I Introduction
To improve diffusion-based generation quality for audio and music, a high-fidelity vocoder is essential. Current methods to do so largely come in two forms: bandwidth extension (expansion) and mono-to-stereo upmixing. Bandwidth extension (BWE) is the process of increasing the frequency range or bandwidth of an audio signal and can be achieved using a source-filter model [16], time-domain [17, 18, 19], or spectral-domain [20, 21, 22, 23, 24] neural network. Mono-to-stereo (M2S) upmixing is the process of converting a single-channel audio signal into spatialized (left and right) channels. M2S can be achieved via decorrelation [25] or a recent parametric stereo method [26] that describes audio as a single-channel plus parameters that approx. characterizes the stereo image [27].
Figure 1: (a) MusicHiFi inference diagram. (b) MusicHiFi unified GAN-training diagram. Blocks from top to bottom represent our vocoder, BWE and M2S. The blue dashed rectangles have shared architectures per cascade stage with different weights.
In this work, we propose MusicHiFi — a new efficient high-fidelity stereophonic vocoder as shown in Fig. 1. Our method employs a GAN-triplet cascade that converts low-resolution single-channel mel-spectrograms (e.g., 22.05 kHz) to stereo high-resolution waveforms (e.g., stereo 44.1 kHz) leveraging vocoder, BWE, and M2S modules, as illustrated in Fig. 0(a). Our method can be seamlessly integrated into low-sampling rate mel-spectrograms music generators, used to enhance the fidelity of a low-resolution audio recording, and/or used to spatialize monophonic music. Compared to recent diffusion-based BWE methods [24] and AR spatialization methods, our GAN-based generators provide much faster GPU inference speed and are easily differentiable. We evaluate our approach via multiple objective and subjective listening tests and find our approach yields equal or better vocoding, BWE, and M2S quality with significantly faster inference speed. Our contributions include:
@@
  • A new fast bandwidth extension method that maintains the low-frequency content in upsampled audio through skip connections, and
@@ II Background
The most relevant recent works to our proposed method include BigVGAN [15] and the Descript audio codec [29]. BigVGAN is a recently proposed vocoder method that has achieved state-of-the-art performance for generating high-fidelity waveforms from mel-spectrograms at speeds significantly faster than real time on a single GPU [15]. The BigVGAN generator employs a neural network architecture composed of a stack of transposed 1D convolutions, each followed by an anti-aliasing multi-periodicity composition (AMP) block which internally uses a Snake activation function [30]. Past work demonstrated AMP blocks are able to generate waveforms with fewer high-frequency artifacts and provide substantial improvements in both objective and subjective assessments. Furthermore, AMP blocks have been found to improve robustness for out-of-distribution vocoding and strong extrapolation ability.
Related to BigVGAN is the Descript audio codec (DAC) neural compression method [31, 32, 29]. DAC uses the SoundStream generator architecture [31] with Snake activations together with an improved GAN-based discriminator architecture, an updated training objective, and a residual vector quantization scheme [31] to achieve state-of-the-art high-fidelity compression. When we look at the discriminator differences, both the BigVGAN and DAC use time-domain multi-period discriminators (MPD) to capture multiple periodic structures as well as spectral-domain discriminators, but DAC replaces the BigVGAN magnitude-only spectral discriminators with a multi-band multi-resolution complex spectrogram discriminators (MMSD) to enhance high-frequency prediction and mitigate aliasing [29]. Both BigVGAN and DAC also leverage a reconstruction loss and an adversarial loss, but DAC includes a codebook loss and updates the reconstruction loss to use multiple mel-bins over multi-scale spectrograms to improve training stability and convergence speed.
@@ III-A Overview
We introduce MusicHiFi, a new vocoding method based on unified triplet-GAN cascade that progressively upsamples audio as shown in Fig. 0(a). Our approach involves three stages, each being modular and useful for different applications. First, low-resolution single-channel mel-spectrograms are transformed into waveforms of the same resolution using a vocoder (MusicHiFi-V). Then, low-resolution waveforms are converted into high-resolution waveforms via our BWE module (MusicHiFi-BWE). Finally, single-channel high-resolution waveforms are upmixed to stereo audio through our M2S module (MusicHiFi-M2S). At each stage, we use an identical generator architecture, discriminator architecture, training objective, and model sizes as shown in Fig. 0(b). The only difference between our three modules are the inputs and outputs and a residual connection for our BWE.
In more detail, for all three of our generator stages, we adopt the BigVGAN transpose 1D convolution + AMP block generator architecture [15] that inputs a mel-spectrogram and outputs audio. For our discriminator architecture, we use the DAC discriminator [29]. For our training objective, we also adopt the DAC reconstruction loss and adversarial loss and remove the codebook learning objective since our focus is high-fidelity audio synthesis instead of neural compression. Our final training objective for our generator ($\mathcal{L}_{G}$) and discriminator ($\mathcal{L}_{D}$) are:
@@
where $D_{k}$ denotes the $k$-th subdiscriminator from MPD or MMSD, $\mathcal{L}_{adv},\mathcal{L}_{rc},\mathcal{L}_{fm}$ represent least-square adversarial loss [33], reconstruction loss and $L_{1}$-based feature matching loss, respectively. Specifically, $\mathcal{L}_{rc}=\sum_{i}||\log{S_{i}}-\log{\hat{S_{i}}}||_{1}$, where $S_{i}$ indicates the $i$-th mel-spectrogram from a list of mel-spectrograms with different fixed resolutions. $L_{fm}$ aims to minimize the distances between real and generated intermediate features from the discriminator layers [13].
@@ III-B MusicHiFi-V
Our vocoder inputs mel-spectrograms computed from low sampling-rate and outputs low-resolution audio waveforms, but otherwise follows our unified generator, discriminator, and training recipe described above. We note that the original BigVGAN training recipe exhibits instabilities and is susceptible to mode collapse when scaled up to larger models [15]. We employ fewer convolution layers in the AMP blocks, extend the input sequence length from 8192 to 16384, increase the convolution channel width of 2048, and use our training strategy, we do not observe mode collapse. Furthermore, this configuration approximately matches the floating point operations per second (FLOPS) of HiFi-GAN [14].
@@ III-C MusicHiFi-BWE
Our BWE module inputs low-resolution audio and outputs high-resolution audio while using our unified generator, discriminator, and training recipe. For our generator architecture, however, we make two small, but significant changes. First, we take the low-resolution audio and compute an intermediate mel-spectrogram representation with half the hop size used for the vocoder to double the sequence length and match full-band waveform output.
Second, we add a residual or skip connection between the input low-resolution audio signal and high-resolution audio output, with a sinc interpolation block in between that performs (2x) upsampling. The residual connection enables our BWE generator to more easily generate low-bandwidth content and allows our BWE generator to focus model capacity on generating high-frequency content. The discriminator also operates on the higher sample rate, full bandwidth audio [20, 21]. Note, during preliminary testing, we experimented not using the residual connection which did not work and applying a low-pass (LP) filter to the generated waveform, which slowed training and did not enhance performance.
@@ IV-A Training details
We train all of our models using an internal dataset of 1800 hours of licensed instrumental music (stereo 44.1 kHz). For training, we randomly crop a sequence of 16,384 samples and then apply module-specific pre-processing. For our vocoder, we use channel averaging and downsampling to 22.05 kHz with STFT settings of a 1024-sample window, 256-sample hop size, and 128-band log-mel spectrogram. For our BWE module, we use the same pre-processing as the vocoder but halve the window and hop size. For our M2S module, we use channel averaging with identical STFT settings as the vocoder. We use the scalar weights $\lambda_{fm}=1$ and $\lambda_{rc}=360$ for the training objective in Eq. (1) and train all modules with a batch size of 45 for 500k steps, and select the optimal checkpoint via the minimum validation multi-resolution STFT distance. Model sizes per stage are approximately 46M params and 55 GFLOPS for one second of audio. All other parameters follow from the open-source implementations of the BigVGAN [15] generator and DAC [29] discriminator.
@@ IV-C Objective evaluation
Dataset Method Mel-Dnormal-$\downarrow$ STFT-Dnormal-$\downarrow$ ViSQOLnormal-$\uparrow$ SI-SDRnormal-$\uparrow$ RTFnormal-$\uparrow$
DSD100 HiFi-GAN [14] 1.09 0.65 4.47 28.62 3488
HiFi-GAN-large [14] 1.06 0.60 4.48 30.29 3409
BigVGAN [15] 0.94 0.41 4.61 27.75 1818
MusicHiFi-V 0.87 0.33 4.67 28.49 1786
FMA HiFi-GAN [14] 1.09 0.64 4.52 32.45 3703
HiFi-GAN-large [14] 1.04 0.57 4.56 31.89 3614
BigVGAN [15] 0.94 0.41 4.62 30.01 1829
MusicHiFi-V 0.87 0.35 4.67 31.57 1807
@@
Dataset Method Mel-Dnormal-$\downarrow$ STFT-Dnormal-$\downarrow$ ViSQOLnormal-$\uparrow$ RTFnormal-$\uparrow$
@@
Dataset Method Mel-Dnormal-$\downarrow$ STFT-Dnormal-$\downarrow$ ViSQOLnormal-$\uparrow$ RTFnormal-$\uparrow$
@@
For objective evaluation, we use 673 clips from FMA-small [37] and the test split of the accompaniment track from the DSD100 test dataset [38]. For both test datasets, every segment has a duration of 30 seconds. For objective evaluation metrics, we use a suite of four metrics including ViSQOL [39], mel distance (Mel-D), multi-resolution STFT distance (STFT-D), and scale-invariant source-to-distortion ratio (SI-SDR) [40]. ViSQOL is a perceptual quality metric that estimates a mean opinion score based on the spectral similarity to the ground truth. The Mel-D and STFT-D measure the spectral distances between the reconstructed and ground-truth audio under a mel and a linear frequency scale, respectively. We also use the SI-SDR metric to measure vocoder quality, but not for BWE and M2S as we found the metric was less reliable for these more multi-modal tasks. We use the real time factor (RTF) metric or time processed over time elapsed on an A100 GPU to measure speed.
@@ IV-D Objective evaluation result
Vocoder objective evaluation results are shown in Table I. We find our proposed method outperforms BigVGAN and HiFi-GAN on Mel-D, STFT-D and ViSQOL results while maintaining a comparable SI-SDR score and lower RTF on both datasets. We also find the RTF of our vocoder method is lower compared to HiFi-GAN, but is still very fast and almost 2000x real-time on an A100.

### v3

@@ Abstract
Diffusion-based audio and music generation models commonly perform generation by constructing an image representation of audio (e.g., a mel-spectrogram) and then convert it to audio using a phase reconstruction model or vocoder. Typical vocoders, however, produce monophonic audio at lower resolutions (e.g., 16-24 kHz), which limits their usefulness. We propose MusicHiFi — an efficient high-fidelity stereophonic vocoder. Our method employs a cascade of three generative adversarial networks (GANs) that convert low-resolution mel-spectrograms to audio, upsamples to high-resolution audio via bandwidth extension, and upmixes to stereophonic audio. Compared to past work, we propose 1) a unified GAN-based generator and discriminator architecture and training procedure for each stage of our cascade, 2) a new fast, near downsampling-compatible bandwidth extension module, and 3) a new fast downmix-compatible mono-to-stereo upmixer that ensures the preservation of monophonic content in the output. We evaluate our approach using objective and subjective listening tests and find our approach yields comparable or better audio quality, better spatialization control, and significantly faster inference speed compared to past work. Sound examples are at https://MusicHiFi.github.io/web/.
@@ I Introduction
To improve diffusion-based generation quality for audio and music, a high-fidelity vocoder is essential. Current methods to do so largely come in two forms: bandwidth extension (expansion) and mono-to-stereo upmixing. Bandwidth extension (BWE) increases the frequency range or bandwidth of an audio signal and can be achieved using a source-filter model [16], time-domain [17, 18, 19], or spectral-domain [20, 21, 22, 23, 24] neural network. Mono-to-stereo (M2S) upmixing converts a single-channel audio signal into spatialized (left and right) channels. M2S can be achieved via decorrelation [25] or a recent parametric stereo method [26] that describes audio as a single-channel plus parameters that approx. characterizes the stereo image [27].
Figure 1: (a) MusicHiFi inference via a cascade of our vocoder, bandwidth extension (BWE), and mono-to-stereo (M2S) modules that use a shared architecture, but with different weights. (b) MusicHiFi GAN-training for our vocoder, BWE and M2S (top-to-bottom), separately. Our unified inference and training scheme enables novel, high-performing BWE and M2S.
We propose MusicHiFi — a new efficient high-fidelity stereophonic vocoder shown in Fig. 1(a). Our method uses a GAN-triplet cascade that converts low-resolution single-channel mel-spectrograms (e.g., 22.05 kHz) to stereo high-resolution waveforms (e.g., stereo 44.1 kHz) via vocoder, BWE, and M2S modules. Our method can be integrated into mel-spectrograms music generators, used to enhance recording fidelity, and/or for spatialization. Compared to diffusion-based BWE methods [24] and AR spatialization, our GAN-based method is much faster and differentiable. We evaluate our approach via objective and subjective listening tests and find our approach yields equal or better vocoding, BWE, and M2S quality with much inference speed. Our contributions are:
@@
  • A new fast BWE method that maintains low-frequency content in upsampled audio via skip connections, and
@@
By creating a unified approach to vocoding, BWE, and M2S, we achieve a significant improvement in audio quality, spatialization, and inference speed compared to past methods.
@@ II Background
The most relevant recent works to our method include BigVGAN [15] and the Descript audio codec (DAC) [29]. BigVGAN is a recently proposed vocoder method that has achieved state-of-the-art performance for generating high-fidelity waveforms from mel-spectrograms at speeds significantly faster than real time on a single GPU [15]. The BigVGAN generator employs a neural network architecture composed of a stack of transposed 1D convolutions, each followed by an anti-aliasing multi-periodicity composition (AMP) block which internally uses a Snake activation function [30]. Past work demonstrated AMP blocks are able to generate waveforms with fewer high-frequency artifacts and provide substantial improvements in both objective and subjective assessments. Furthermore, AMP blocks have been found to improve robustness for out-of-distribution vocoding and strong extrapolation ability.
DAC is a neural-based codec that follows the SoundStream generator [31] architecture with Snake activations, an improved GAN-based discriminator, an enhanced reconstruction training objective, and a residual vector quantization scheme that achieves state-of-the-art compression. When we look at the discriminator differences, both the BigVGAN and DAC use time-domain multi-period discriminators (MPD) to capture multiple periodic structures as well as spectral-domain discriminators, but DAC replaces the BigVGAN magnitude-only spectral discriminators with a multi-band multi-resolution complex spectrogram discriminators (MMSD) to enhance high-frequency prediction and mitigate aliasing [29]. Compared to the reconstruction loss proposed by BigVGAN, DAC uses multiple mel-bins over multi-scale spectrograms to improve training stability and convergence speed.
@@ III-A Overview
We introduce MusicHiFi, a new vocoding method based on unified triplet-GAN cascade that progressively upsamples audio as shown in Fig. 1(a). Our approach involves three stages, each being modular and useful for different applications. First, single-channel mel-spectrograms of low sampling rate are transformed into waveforms of the same sampling rate using a vocoder (MusicHiFi-V). Then, waveforms of low sampling rate are converted into waveforms with full-bandwidth via our BWE module (MusicHiFi-BWE). Finally, the single-channel waveforms are upmixed to stereo audio through our M2S module (MusicHiFi-M2S). At each stage, we use an identical generator architecture, discriminator architecture, training objective as shown in Fig. 1(b). The main differences between the three modules are the varying input and output dimensions, and the inclusion of an additional skip connection from the input in the BWE module.
In more detail, for all three of our generator stages, we adopt the BigVGAN transpose 1D convolution + AMP block generator architecture [15] that inputs a mel-spectrogram and outputs audio. For our discriminator architecture, we use the DAC discriminator [29]. For our training objective, we also adopt the DAC reconstruction loss and adversarial loss and remove the codebook learning objective since our focus is high-fidelity audio synthesis. Our final training objective for our generator ($\mathcal{L}_{G}$) and discriminator ($\mathcal{L}_{D}$) are:
@@
where we apply $K$ discriminators, $D_{k}$ denotes the $k$-th subdiscriminator from MPD or MMSD, $\mathcal{L}_{adv},\mathcal{L}_{rc},\mathcal{L}_{fm}$ represent least-square adversarial loss [32], reconstruction loss and $L_{1}$-based feature matching loss, respectively and $\lambda_{rc}$ and $\lambda_{fm}$ represent corresponding loss weighting term. Specifically, $\mathcal{L}_{rc}=\sum_{i}||\log{S_{i}}-\log{\hat{S_{i}}}||_{1}$, where $S_{i}$ indicates the $i$-th mel-spectrogram from a list of mel-spectrograms with different fixed resolutions. $L_{fm}$ aims to minimize the distances between real and generated features from discriminator layers [13].
@@ III-B MusicHiFi-V
Our vocoder inputs mel-spectrograms of low sampling-rate and outputs audio waveforms with the same sampling rate, and follows our unified generator, discriminator, and training recipe described above. We note that the original BigVGAN training recipe exhibits instabilities and is susceptible to mode collapse when scaled up to larger models [15]. To enhance performance while stabilizing training, we double the input length of the audio samples originally defined in BigVGAN, reduce the number of convolution layers in the AMP blocks, increase the convolution channel width to 2048. Furthermore, this configuration approximately matches the floating point operations per second (FLOPS) of HiFi-GAN [14].
@@ III-C MusicHiFi-BWE
Our BWE module takes audio with a low sampling rate as input and outputs full-band audio and follows our unified generator and discriminator architectural design, and training objectives. For our generator architecture, however, we make two small, but significant changes. First, we compute an intermediate mel-spectrogram representation for the input audio with half the hop size used for the vocoder to double the sequence length and match full-bandwidth waveform output.
Second, we add a residual or skip connection between the input audio signal of low sampling rate and the full-bandwidth audio output, with a sinc interpolation block in between that performs 2x upsampling. The residual connection enables our BWE generator to more easily generate low-bandwidth content and allows our BWE generator to focus model capacity on generating high-frequency content. The discriminator also operates on the higher sample rate, full bandwidth audio [20, 21]. Note, during preliminary testing, we experimented not using the residual connection which did not work and applying a low-pass (LP) filter to the generated waveform, which slowed training and did not enhance performance.
@@ IV-A Training details
We train all of our models using an internal dataset of 1800 hours of licensed instrumental music (stereo 44.1 kHz). For training, we randomly crop a sequence of 16,384 samples and then apply module-specific pre-processing. For our vocoder, we use channel averaging and downsampling to 22.05 kHz with STFT settings of a 1024-sample window, 256-sample hop size, and 128-band log-mel spectrogram. For our BWE module, we use the same pre-processing as the vocoder but halve the window and hop size. For our M2S module, we use channel averaging with identical STFT settings as the vocoder. We use the scalar weights $\lambda_{fm}=1$ and $\lambda_{rc}=360$ for the training objective in Eq. (1) and train all modules with a batch size of 45 for 500k steps, and select the optimal checkpoint via the minimum validation multi-resolution STFT distance (STFT-D). Model sizes per stage are approximately 46M params and 55 GFLOPS for one second of audio. All other parameters follow from the open-source implementations of the BigVGAN [15] generator and DAC [29] discriminator.
@@ IV-C Objective evaluation
[…]
DSD100 HiFi-GAN [14] 1.09 0.65 4.47 3488
HiFi-GAN-large [14] 1.06 0.60 4.48 3409
BigVGAN [15] 0.94 0.41 4.61 1818
MusicHiFi-V 0.87 0.33 4.67 1786
FMA HiFi-GAN [14] 1.09 0.64 4.52 3703
HiFi-GAN-large [14] 1.04 0.57 4.56 3614
BigVGAN [15] 0.94 0.41 4.62 1829
MusicHiFi-V 0.87 0.35 4.67 1807
@@
[…]
@@
[…]
@@
For objective evaluation, we use 673 clips from FMA-small [36] and the test split of the accompaniment track from the DSD100 test dataset [37]. For both test datasets, every segment has a duration of 30 seconds. For objective evaluation metrics, we use a suite of four metrics including ViSQOL [38], mel distance (Mel-D) and STFT-D. ViSQOL is a perceptual quality metric that estimates a mean opinion score based on the spectral similarity to the ground truth. The Mel-D and STFT-D measure the spectral distances between the reconstructed and ground-truth audio under a mel and a linear frequency scale, respectively. We use the real time factor (RTF) metric or time processed over time elapsed on an A100 GPU to measure speed.
@@ IV-D Objective evaluation result
Vocoder objective evaluation results are shown in Table I. We find our proposed method outperforms BigVGAN and HiFi-GAN on Mel-D, STFT-D and ViSQOL results while maintaining a lower RTF on both datasets. We also find the RTF of our vocoder method is lower compared to HiFi-GAN, but is still very fast and almost 2000x real-time on an A100.